OpenStudy (anonymous):

Evaluate the expression when x = 7 and y = 2. 2(x + y)2

OpenStudy (anonymous):

please help

OpenStudy (bibby):

\(2(x+y)^2\)?

OpenStudy (anonymous):

Yes

OpenStudy (bibby):

PEDMAS/PEMDAS 2(7+2)^2 start with the stuff inside the parentheses

OpenStudy (anonymous):

K that would be 9 right?

OpenStudy (bibby):

yeah. what's next? parentheses exponents division/mult add/sub

OpenStudy (anonymous):

then exponents

OpenStudy (bibby):

so what is 2(9)^2?

OpenStudy (anonymous):

162?

OpenStudy (bibby):

2*9^2=2*81=162 yea

OpenStudy (texaschic101):

2(x + y)^2 = 2(7 + 2)^2 = 2(9^2) = 2(81) = 162

OpenStudy (anonymous):

K so..... the answer to my question would be 162 ?

OpenStudy (bibby):

at least 2 people have said yes, 3 if you count you but you don't know anything

OpenStudy (anonymous):

That not nice

OpenStudy (bibby):

it's a joke, soz

OpenStudy (bibby):

you're right

OpenStudy (anonymous):

I know how to do Order Of Operations I just did not know this specific question

OpenStudy (anonymous):

K

OpenStudy (anonymous):

K thanks so much guys